40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 270 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  3310   Wed Jul 28 14:34:29 2010 channaUpdateComputersinstallation on allegra

I have done the following on allegra and rosalba:

[root@allegra caltech]# yum install glade2

On rosalba the matplotlib was out of date with respect to allegra.  I have no idea how the version 0.98 on allegra got there, but I left it.  However I updated rosalba to the epel version

  1 yum remove python-numpy
  2 yum install python-matplotlib numpy scipy --enablerepo=epel --disablerepo=rpmforge

 

This is all to support the LIGO data listener which now has a shortcut on rosalba and allegra's desktop.  It seems to work for (live mode) right now.
 

 

  3309   Wed Jul 28 13:06:47 2010 RazibUpdatePhase Camera 

Attached are some calculation that I did previously for the phasecamera setup. This shows the nature of the beat signal that we are measuring.

I am also trying to characterize the noise source of the camera also. Following images shows the mean dark noise (with no light on the camera) and the standard deviation for 100 snaps at an exposure time of 500 µs.

mean_100_snaps.pngstd_100_snaps.png

My target now is to measure the response gain of each pixel and how they vary over intensity. I already have a simplified setup on the table and will work on it today. Details will follow at the end of the day.

Attachment 3: phase_cam_calc.pdf
phase_cam_calc.pdf phase_cam_calc.pdf phase_cam_calc.pdf
  3308   Wed Jul 28 12:53:32 2010 channaUpdateComputersnds data listener

For the sake of writing it down: /cvs/cds/caltech/apps/linux64/rockNDS

  3307   Wed Jul 28 12:31:00 2010 GopalUpdateWIKI-40M Update7.21.10-7.28.10 Weekly Update

Summary of this week's activities:

7/21: Frequency Domain Analysis of rectangular bar; discussed with Koji how to convert complex eigenfrequencies into phase factors.

7/23: Created Wiki page about FDA; Journal Club

7/26: Recreated Stack_1234.mph due to boundary value issues; FDA for 1,2,3,4,5 Hz

7/27: Discovered MC2 logbooks for later design; ran the complete x-translational FDA for Stack_1234.mph

7/28: Finished y-translational FDA (posted previously); "Tapered Cantilever" COMSOL tutorial for gravity-load analysis.

  3306   Wed Jul 28 12:16:03 2010 GopalUpdateSEIBode Magnitude Plot and Concerns

Quote:

1) Gravity has to be included because the inverted pendulum effect changes the resonant frequencies. The deflection from gravity is tiny but the change in the dynamics is not. The results are not accurate without it. The z-direction probably is unaffected by gravity, but the tilt modes really feel it.

2) You should try a better meshing. Right now COMSOL is calculating a lot of strain/stress in the steel plates. For our purposes, we can imagine that the steel is infinitely stiff. There are options in COMSOL to change the meshing density in the different materials - as we can see from your previous plots, all the action is in the rubber.

3) I don't think the mesh density directly limits the upper measurement frequency. When you redo the swept-sine using the matlab scripting, use a logarithmic frequency grid like we usually do for the Bode plots. The measurement axis should go from 0.1 - 30 Hz and have ~100 points.

In any case, the whole thing looks promising: we've got real solid models and we're on the merge of being able to duplicate numerically the Dugolini-Vass-Weinstein measurements.

I made some progress on a couple issues:

1) I figured out how to create log-transfer function plots directly in COMSOL, which eliminates the hassle of toggling between programs.

2) Instead of plotting maximum displacement, which could lead to inconsistencies, I've started using point displacement, standardizing to the center of the top surface.

3) I discovered that the displacement can be measured as a field vector, so the minor couplings between each translational direction (due to the asymmetry in the original designs) can be easily ignored. 

Bode_Disp_MC1_MC3_y.png

  3305   Wed Jul 28 12:09:06 2010 Sharmila,KatharineUpdateWIKI-40M UpdateMaglev

We have modeled our maglev setup in simulink but we have a few corrections to make since the system goes into undamped oscillations for an impulse in the input.

We have made a stable mount for the system and started to work on the 2X2 system using this mount. We have to figure out a way to match the magnets with the gain. We have attached the simulink block.

Picture_1.png

  3304   Wed Jul 28 01:05:44 2010 ranaUpdateSEIBode Magnitude Plot and Concerns

1) Gravity has to be included because the inverted pendulum effect changes the resonant frequencies. The deflection from gravity is tiny but the change in the dynamics is not. The results are not accurate without it. The z-direction probably is unaffected by gravity, but the tilt modes really feel it.

2) You should try a better meshing. Right now COMSOL is calculating a lot of strain/stress in the steel plates. For our purposes, we can imagine that the steel is infinitely stiff. There are options in COMSOL to change the meshing density in the different materials - as we can see from your previous plots, all the action is in the rubber.

3) I don't think the mesh density directly limits the upper measurement frequency. When you redo the swept-sine using the matlab scripting, use a logarithmic frequency grid like we usually do for the Bode plots. The measurement axis should go from 0.1 - 30 Hz and have ~100 points.

In any case, the whole thing looks promising: we've got real solid models and we're on the merge of being able to duplicate numerically the Dugolini-Vass-Weinstein measurements.

  3303   Tue Jul 27 23:46:45 2010 JenneUpdateSUSQ measurements of 2 TTs

[Koji, Jenne]

We took measurements of the Q of all the modes that we could think of for TT#4, and then repeated several of the same measurements for TT#2.  We noticed that when we took off the backplane and then replaced it on TT#4, the pitch pointing had changed, so we had to repeat the balancing procedure by slightly shifting the position of the wire clamps relative to the mirror holder. No fun. We decided to quit removing the backplanes. 

The main conclusion of this evening's measurements of TT#4 is that everything looks very close to the design ideas.  Good work team!

TT#4:

'Free Swinging' values (just for interest)

Vert, no damping:   Q = 31.4

Pitch, no damping (ECD backplane removed): Q = ~700

Yaw, no ECDs: Q = ~900

Pos, no ECDs (no measurement) - we had already put the backplane back on, and didn't want to take it off again.

 

Damped Values:

Vert, with damping: Q = 14.3

Pitch, with ECDs: overdamped, so Q < 1/2

Yaw, with ECDs: Q = 2.3

Pos, with ECDs: Q = 1.4

Side, with ECDs: Q = 1.9

 

We also measured the resonant frequency of each of the ECDs for this TT (since we had the backplane removed anyway...)

ECD UL: 10.05Hz

ECD UR: 10.15Hz

ECD LL: 10.21Hz

ECD LR: 10.21Hz

 

TT#2:

Yaw, with ECDs: Q = 7.0

Pitch, with ECDs: overdamped, so Q < 1/2

Vert: Problematic.  No damping, f = 25.9Hz, Q = 36.  With rubber dampers, f = 20.0Hz, Q = 42.   Yes, you read that right.  The frequency is lower, and the Q is higher *with* the damping.  Perhaps our brains are fried.  Perhaps we've discovered new, inconsistent physics (awfully unlikely....). We'll revisit this again tomorrow to figure out what mistake we're making.

  3302   Tue Jul 27 21:38:21 2010 kiwamuUpdateSUSinstallation of in-vac optics

[Alberto and Kiwamu] 

We put the PRM back approximately on the right place.

Also we installed the pick off mirrors and the PZT mirror.

Since the main beam after the MMT still has not been well aligned , we put those optics approximately on the right place. A fine alignment of those will be performed later

The offsets of the PRM OSEMs are still kind of okay.

The next things we have to do are

(1) installation of a tip-tilt for the SRC, (2) alignement of those optics by using the main laser and (3) installation of the green optics.

 


what we did

 1. put the PRM back to the designed place.

     - After this, we released the PRM from the earthquake stops and turned on the damping servo.

      - Now the earthquake stops are at a distance of approximately 1mm from the PRM. These separation distances were tuned by counting the turn number when we screwed them off.

 2. leveled the table

 3. adjusted the separation distances from the PRM to the OSEMs.

    - The table below summarize the current OSEM offsets. LL may still need to be adjusted.

  Max./2 [V]

measured offsets [V] 

after adjustment 

discrepancy [%]
UL 0.874 0.928 + 6.2 %
UR 0.848 0.777 - 8.4 %
LL 1.038 0.796 - 23 %
LR 0.967 0.845 -12 %
SD 0.840 0.745 - 11%

 4. put the PZT mirror on the right place.

      -  This PZT mirror is going to be used for beam steering after the MMT.

 5. put the pick off mirror and its associated optics.

     -  This pick off mirror provides with the beam eventually going to IP_ANG and IP_POS.

 


current status

The table below shows the current status of the installed optics.

Red letters represent the incomplete states which still need further adjustment.

Blue letters represent the complete status which don't need any further adjustment.

 

name on the drawing

(see the wiki ) 

status
BS tower BS well aligned by a caliper
PRM tower PRM approximately aligned. LL OSEM
TIp-TIlt PR3 approximately aligned
Pick off Window (wedged) IPPO

wedge is correctly set (fat part is on the left).

 approximately aligned

steering mirror IPPOSSM1 mirror is correctly flipped. approximately aligned
steering mirror IPANGSM1 mirror is correctly flipped. approximately aligned
steering mirror IPANGSM2 approximately aligned
pzt mirror SM2 approximately aligned

 

Attachment 1: DSC_2284.JPG
DSC_2284.JPG
Attachment 2: DSC_2291.JPG
DSC_2291.JPG
  3301   Tue Jul 27 18:42:57 2010 GopalUpdateOptic StacksBode Magnitude Plot and Concerns

I completed the frequency domain analysis mentioned previously in the x-direction. Although I ran it from 1-10 Hz, with 0.1-Hz increments, COMSOL was unable to complete the task past 7 Hz because the relative error was beyond the relative tolerance. To solve this issue, I'd have to rerun the simulation with a finer mesh, an unfavorable option because of the already-extensive run times. The Bode magnitude plot from this simulation is attached:

Bode_Mag_MC1_MC3.png

 

This simulation raises some questions about the feasibility of this method:

 

1) Do we have the computing power necessary?

 

I already moved my work from my personal Mac Pro to Kallo (4 GB --> 12 GB RAM difference). Now, instead of crashing the program constantly, I typically wait a half hour for a standard run of the model. Preferably, I could move my work to Megatron or some other workhorse-computer... but I also know that many of the big boys are already being strained as is.

 

2) Is it possible to take measurements through Matlab?

 

This way, I could write a script to instruct COMSOL and just run a few tests at a time overnight. Also, I wouldn't have to sit and record measurements manually as I've done here. The benefits of such an improvement warrant further attention. I'll work on this option next.

 

3) Up until what frequency do we need to model?

 

As I've shown, normal meshing yields data up to 7 Hz. Is this enough? Do we need more data? Certainly not less, I'm quite sure. We need to keep in mind that as frequency range increases, run times increase exponentially.

 

4) How do we incorporate gravity into the equation?

 

Gravity will produce a bit of extra force in the non-restoring direction for off-axis deviations, slightly decreasing the expected frequency. Whether or not this is an important effect is questionable, since the deviations are typically on the order of a micron, which is orders of magnitude smaller than the initial displacement I'm using on the base. I've decided to ignore this complication for now.

 

 

  3300   Tue Jul 27 16:33:50 2010 KojiSummaryGeneralHigh school students tour

Jenne made the 40m tour for the annual visit of 30-40 students.

c.f.

Tour 2009 http://nodus.ligo.caltech.edu:8080/40m/1717

Tour 2008 http://nodus.ligo.caltech.edu:8080/40m/737

 

Attachment 1: IMG_2657.jpg
IMG_2657.jpg
  3299   Tue Jul 27 16:03:36 2010 ranaSummaryDAQDAQ timing test

Quote:

Since we now have a good measurement of the phase noise of the Rb clock Marconi locked to the Rb clock, I wanted to use that to check out the old DAQ system:

I used Megan's phase noise setup - Marconi #2 is putting out 11000013 Hz at 13 dBm into the ZP-3MH mixer. Marconi #1 is putting out 3 dBm at 11000000 Hz into the RF input.

The output goes through a 50 Ohm load and then a Mini-Circuits BNC LP filter (either 2 or 5 MHz). Then an SR560 set for low noise, G = 5, AC coupling, 1-pole LP @ 1 kHz.

This SR560 output goes into the channel C1:IOO-MC_DRUM1 (which is sampled at 16384 Hz with ICS-110B after the usual Sander Liu AA chassis containing the INA134s).

 This is the 0.3 mHz BW spectrum of this test - as you can see the apparent linewidth (assuming the width is all caused by the DAQ jitter) is comparable to the BW and therefore not resolved.

Basically, the Hanning window function is not sharp enough to do this test and so I will do it offline in Matlab.

Attachment 1: Untitled.png
Untitled.png
  3298   Tue Jul 27 12:02:31 2010 GopalUpdateOptic StacksPreliminary Transfer Function Measurements on MC1/MC3

I have successfully completed a preliminary transfer function measurement test on the MC1/MC3 stack in COMSOL. Using the measurement scheme described on the Wiki, I initialized a 1 N/m^2 sinusoidal perturbation on the bottom of the stack and measured the maximum displacement of the top layer. This preliminary test just calculated the responses to 1-,2-,3-,4-, and 5-Hz drives along the x-axis (pictures attached).

Currently, I am rerunning the same test but from 1-10 Hz with 0.1-Hz steps. When both x- and y-axis responses have been plotted, I can move on to repeating this entire process on the MC2 stack.

Attachment 1: MC1_MC3_FDA_1.png
MC1_MC3_FDA_1.png
Attachment 2: MC1_MC3_FDA_2.png
MC1_MC3_FDA_2.png
Attachment 3: MC1_MC3_FDA_3.png
MC1_MC3_FDA_3.png
Attachment 4: MC1_MC3_FDA_4.png
MC1_MC3_FDA_4.png
Attachment 5: MC1_MC3_FDA_5.png
MC1_MC3_FDA_5.png
  3297   Tue Jul 27 11:43:24 2010 steveUpdatePEMair quality is bad today

The lab is at 30,000 and Pasadena air is at 1.1 e+6 particles /cf min of 0.5 micron.

Attachment 1: pemtoday.jpg
pemtoday.jpg
  3296   Tue Jul 27 11:24:53 2010 josephbHowToComputer Scripts / Programskilldataviewer script

I placed a script for killing all instances of the dataviewer program on the current computer in /cvs/cds/caltech/scripts/general/.  Its called killdataviewer.  This is intended to get rid of a bunch of zombie dataviewer processes quickly.  These processes get into this bad state when the dataviewer program is closed in any way other than the graphical menu File -> Exit option.

Its contents are very simple:

#/bin/bash

kill `ps -ef | grep dataviewer | grep -v grep | grep -v killdataviewer | awk '{print $2}'`

  3295   Mon Jul 26 20:30:35 2010 JenneUpdateSUS2 Tip Tilts suspended and balanced

[Koji, Jenne]

We were on Team Cleanroom, while Kiwamu and Alberto were on Team Chamber.  Team Cleanroom suspended and balanced 2 Tip Tilts this afternoon.

One of the TTs that was suspended today is the one which was broken on Friday (see elog 3278).  We resuspended it using the regular 0.0036" diameter wire (91um).  We balanced it using the HeNe oplev, and then set it aside.  This TT has serial number 2.

We noticed that, like the previous 2 TT suspensions (this one before it was broken, and the one actually installed in the BS chamber on Friday, which is #3), there seems to be a little bit of hysteresis in the pointing.  The difference comes if we poke the top of the mirror holder and observe the place the reflected beam spot comes to rest at, and if we poke the bottom of the mirror holder.  The beam spot stays a little higher when we poke the top vs. when we poke the bottom. 

To combat this, we tried suspending our second TT of the day (the one that Kyung Ha and I had half finished) using thinner wire for the mirror holder.  We used the 0.0017" diameter wire (43um) that is used for the SOSes.  Unfortunately, it still seems like there is a similar hysteresis.  The thin-wire TT has serial number 4.

While working on TT4, we recalled that we have to include rubber dampers for the vertical blade springs.  Oooops!  We used some of the leftover #4-40 screws with viton tips that Zach and Mott had made for Earthquake stops to damp the vertical resonance of the blades.  We measured the Q factor by flicking the blades up or down.  We changed the oplev setup to be a shadow sensor setup, and watched the ringdown of the vertical mode on the 'scope.  We counted #cycles/time = frequency, and the t(1/2) time for the exponential ringdown to calculate the Q.  For the shadow sensor, we positioned the QPD in line with the initial HeNe beam, and placed the edge of the mirror holder clamp partially in the beam, so the beam was partly occluded.  When the mirror shook up and down, more or less of the beam was blocked, and we could see this power fluctuation on the 'scope.

Using the formula Q = pi  f0 T1/2 / ln(2) = 4.53 f0 T1/2, where T1/2 is the the time it takes for the amplitude to decay by half, we measured a Q of 31 for the vertical mode with no damping, and a Q of 14 with damping.  Koji confirmed the calculation and put it into wiki.

We need to go through the other TTs that have been assembled and give them their rubber dampers.

 

  3294   Mon Jul 26 20:12:18 2010 kiwamuUpdateSUSOSEMs on PRM

 [Alberto and Kiwamu]

We installed the OSEMs to the new PRM.

As I wrote down on the elog (see here)  today's mission was to install the OSEMs to the PRM.

After putting them on the tower we adjusted the readout offsets by sliding the OSEMs so that they can stay in the linear sensing ranges. 

Now all of the OSEMs have almost good separation distances from the PRM.

In the attached picture you can see the OSEMs installed on the PRM tower ( middle: PRM tower, left: BS tower)


(what we did)

 1. moved the PRM tower close to the door so that we could easily access the PRM.

 2. leveled the table by putting some weights and confirmed the level by a  bubble level tool.

     - We must level the table every time when we set / adjust any OSEMs,  otherwise the readout voltages of  the OSEMs vary every time due to the tilted table.

 3. released the PRM by loosing the earthquake stops

 4. put the OSEMs with approximately right separation distances from the PRM.

      -  At this phase we can see the readout of the OSEMs, which were oscillating freely because we still didn't enable the damping.

        -  The OSEM positions were checked by looking at useful notes on the wiki (see here).

 5. turned on the damping servo of the OSEMs

       - Without changing any gains, it worked well. 

      - Then we could see stable readouts of the OSEMs which didn't show any oscillations in turn because of the damping.

 6. checked the level of the table again

 7. set each of the OSEM readouts to the half of its maximum value by sliding their positions slightly.

      - The readout offsets were at almost the half value within +/- 100 mV accuracy (this was the best accuracy we could adjust by our hands)

 8. screwed down the earthquake stops to lock the PRM.

      - Now the damping is off.

 9. closed the door

 


(to be done)

 *  Putting the PRM tower back to the designed place

 *  Installation of the pick off mirror

 *  Arrangement of the PZT mirror

Attachment 1: DSC_2279.JPG
DSC_2279.JPG
  3293   Mon Jul 26 14:24:46 2010 josephb, kiwamuUpdateCDSRFM test take 1

Kiwamu and I strung a temporary RFM fiber from the c1iscex machine (in the new 1X9 rack) to the c1sus machine (in the new 1X4 rack).  This was connected into the respective RFM cards.  Once we put the fiber in correctly, the status lights came on the RFM card, which is a good sign.  This did not go through the RFM bypass, and did not interfere with any other RFM connections.

We created a simple model to test the RFM card, which basically was 4 RFM memory locations passing back and forth between 2 filters on each machine.  These models were called c1rf0 (on c1sus) and c1rf1 (on c1iscex).  We added 4 entries to the /cvs/cds/caltech/chans/ipc/C1.ipc file corresponding to the 4 RFM memory locations, set their ipcType=RFM and set the ipcRate to 65536.  The ipcNum were set from 0 to 3. The models ran, however, the data we were trying to pass over the RFM card did not seem to be being passed.  Currently trying to contact Alex via e-mail to get debugging advice, and confirm the ipc file is setup correctly.

  3292   Mon Jul 26 12:31:36 2010 kiwamuUpdateCDSfront end machine for the X end

A brief report about the new front end machine "C1ISCEX" installed on the X end (old Y end).

Still the DAC is not working.

 

- Timing card

It's working correctly.

The 1PPS clock signal is supplied by the vertex clock distributer via a 40m long fiber.

 

- ADC

All 16 channels are working well.

We can see the signals in the medm screen while injecting some signals to the ADC by using a function generator.

 

-DAC

All 16 channels do NOT work.

We can not see any signals at the DAC SCSI cable while digitally injecting a signal on the medm screen.

  3291   Mon Jul 26 11:15:23 2010 GopalHowToCOMSOL TipsPictures from Transfer Function Tutorial on the Wiki

The attached pictures give a brief overview of my transfer function measurement procedure in COMSOL. For more details, please see the Wiki.

Screen_shot_2010-07-23_at_2.57.14_PM.png

Screen_shot_2010-07-23_at_2.57.38_PM.png

Screen_shot_2010-07-23_at_2.57.45_PM.png

Screen_shot_2010-07-23_at_2.58.05_PM.png

Screen_shot_2010-07-23_at_2.58.18_PM.png

Screen_shot_2010-07-23_at_2.59.02_PM.png

Screen_shot_2010-07-23_at_3.00.37_PM.png

  3290   Mon Jul 26 10:04:24 2010 steveUpdateGeneralInitial Crane Inspection reveals flaws: wiring, oil and imbalance

Quote:

The guy from KroneCrane (sp?) came today and started the crane inspection on the X End Crane. There were issues with our crane so he's going to resume on Monday. We turned off the MOPA fur the duration of the inspection.

  1. None of our cranes have oil in the gearbox and it seems that they never did since they have never been maintained. Sloppy installation job. The crane oiling guy is going to come in on Monday.
  2. They tried to test the X-End crane with 2500 lbs. (its a 1 ton crane). This tripped the thermal overload on the crane as intended with this test. Unfortunately, the thermal overload switch disabled the 'goes down' circuit instead of the 'goes up' circuit as it should. We double checked the wiring diagram to confirm our hypothesis. Seems the X-End crane was wired up incorrectly in the first place 16 years ago. We'll have to get this fixed.

The plan is that they will bring enough weight to test it at slightly over the rating (1 Ton + 10 %) and we'll retry the certification after the oiling on Monday.

 The south end crane has one more flaw. The wall cantilever is imbalanced: meaning it wants to rotate south ward, because its axis is off.

This effects the rope winding on the drum as it is shown on Atm2

Atm1 is showing Jay Swar of KoneCrane and the two 1250 lbs load that was used for the test. Overloading the crane at 125% is general practice at load testing.

It was good to see that the load brakes were working well at 2500 lbs. Finally we found a good service company! and thanks for Rana and Alberto

for coming in on Saturday.

Attachment 1: 2500.JPG
2500.JPG
Attachment 2: sedrum.JPG
sedrum.JPG
  3289   Mon Jul 26 10:02:42 2010 josephbUpdateCDSRerouted RFM around c1lsc, took RFM card out of c1lsc

If you're refering to just the medm screen,  those can be restored from the SVN.  As we're moving to a new directory structure, starting with /opt/rtcds/caltech/c1/, the old LSC screens can all be put back in the /cvs/cds/caltech/medm/c1/lsc directory if desired.

The slow lsc aux crate, c1iscaux2, is still working, and those channels are still available.  I confirmed that one was still updating. As a quick test, I went to the SVN and pulled out the C1LSC_RFADJUST.adl file, renamed it to C1LSC_RFadjust.adl and placed it in /cvs/cds/caltech/medm/c1/lsc/, and checked it linked properly from the C1IOO_ModeCleaner.adl file.  I haven't touched the modulation depths, as I didn't want to mess with the mode cleaner, but if I get an OK, we can test that today and confirm that modulation depth control is still working.

Quote:

 I just realized that an unfortunate casualty of this LSC work was the deletion of the slow controls for the LSC which we still use (some sort of AUX processor). For example, the modulation

depth slider for the MC is now in an unknown state.

 

  3288   Mon Jul 26 08:51:31 2010 steveUpdatePEM lowering AC setpoint by 2F makes PSL happy

Quote:

The control room temp is warmer than usual. The heat exchanger Office Pro 18 set point was lowered from 70 to 68F yesterday.

The MOPA headtemp is higher also. The Neslab chiller bath temp peaks around 21.6 C daily. This should be rock solid 20.00 C

It did not have any effect.

Now, I have just lowered the thermostat setting of room 101 from 73 to71F  I  hope Koji can take this.

 Little change in the AC set temp can make wonders. Neslab chiller bath temp 19.99C is back to normal and daily variation of PSL are much better.

Attachment 1: htempRTemp.jpg
htempRTemp.jpg
  3287   Sun Jul 25 18:47:23 2010 AlbertoUpdateSVNOptickle 40mUpgrade model updated to include short cavity length corrections

I uploaded an updated optickle model of the upgrade to the SVN directory with the optickle models (here).

  3286   Sat Jul 24 14:27:36 2010 ranaUpdateElectronicsFSS Oscilaltor Phase Noise Measurement

I reconnected the RF signal to the FSS and to the FSS' EOM so that we could lock the refcav again.

I then started a 3 sec. period trianglewave on the AOM drive amplitude to see if there is a direct coupling from RIN to Frequency. Ideally we will be able to measure this by looking at the RCTRANS and the FSS-FAST.

  3285   Sat Jul 24 14:03:19 2010 AlbertoUpdateElectronicsFSS Oscilaltor Phase Noise Measurement

[Rana, Alberto]

Today we measured the phase noise of the oscillator used for the FSS.

The source is a Wenzel crystal at about 21.5MHz that Peter Kalmus built some time ago.

We basically used the same technique that Frank and Megan have been using lately to measure the Marconi's phase noise.

Today we just did a quick measurement but today next week we are going to repeat it more carefully.

Attached is a plot that shows the measurement calibrated for a UGF at about 60 Hz. The noise is compared to that specified by Wenzel for their crystal.

The noise is bigger than that of the MArconi alone locked to the Rubidium standard (see elog entry). We don't know the reason for sure yet.

We'll get back to this problem next week.

Attachment 1: FSScrystalPhaseNoiseHigherGain.pdf
FSScrystalPhaseNoiseHigherGain.pdf
  3284   Sat Jul 24 13:13:41 2010 rana, steve, albertoUpdateGeneralInitial Crane Inspection reveals flaws: wiring and oil

The guy from KroneCrane (sp?) came today and started the crane inspection on the X End Crane. There were issues with our crane so he's going to resume on Monday. We turned off the MOPA fur the duration of the inspection.

  1. None of our cranes have oil in the gearbox and it seems that they never did since they have never been maintained. Sloppy installation job. The crane oiling guy is going to come in on Monday.
  2. They tried to test the X-End crane with 2500 lbs. (its a 1 ton crane). This tripped the thermal overload on the crane as intended with this test. Unfortunately, the thermal overload switch disabled the 'goes down' circuit instead of the 'goes up' circuit as it should. We double checked the wiring diagram to confirm our hypothesis. Seems the X-End crane was wired up incorrectly in the first place 16 years ago. We'll have to get this fixed.

The plan is that they will bring enough weight to test it at slightly over the rating (1 Ton + 10 %) and we'll retry the certification after the oiling on Monday.

  3283   Fri Jul 23 21:35:48 2010 ranaUpdateCDSRerouted RFM around c1lsc, took RFM card out of c1lsc

 I just realized that an unfortunate casualty of this LSC work was the deletion of the slow controls for the LSC which we still use (some sort of AUX processor). For example, the modulation

depth slider for the MC is now in an unknown state.

  3282   Fri Jul 23 21:14:29 2010 RanaUpdatePSLReference Cavity Insulation

I wrapped another ~3 layers onto there. It occurs to me now that we could just get some 2mm thick copper plates made to fit over the stainless steel can.

They don't have to completely cover it, just mostly. I also took the copper circles that Steve had made and marked them with the correct beam height

and put them on Steve's desk. We need a 1" dia. hole cut into these on Monday.

To compensate for the cooling during my work, I've set the heater for max heating for 1 hour and then to engage the temperature servo.

I also noticed the HEPA VARIAC on the PSL was set to 100. Please set it back to 20 after completing your PSL work so that it doesn't disturb the RC temperature..

  3281   Fri Jul 23 18:55:25 2010 kiwamuUpdateSUSPRM and TT installed

I updated the last entry.

  3280   Fri Jul 23 16:02:16 2010 RanaUpdatePSLReference Cavity Insulation

This is the trend so far with the copper foil wrapping. According to Megan's calculation, we need ~1 mm of foil and the thickness of each layer is 0.002" (1/20th of a mm), so we need ~20 layers. We have ~5 layers so far.

As you can see the out-of-loop temperature sensor (RCTEMP) is much better than before. We need another week to tell how well the frequency is doing -

the recent spate of power cycles / reboots of the PSL have interrupted the trend smoothness so far.

Attachment 1: Untitled.png
Untitled.png
  3279   Fri Jul 23 16:00:35 2010 steveUpdateSAFETYcrane load test tomorrow

All 3 cranes will be load tested at 1 ton tomorrow morning between 9am and 2pm

Do not come to the 40m lab during this period. We may disturb your experiment.

Please prepare your touchy set ups to take this test.

  3278   Fri Jul 23 15:54:38 2010 kiwamuUpdateSUSPRM and TT installed

[ Jenne, Koji and Kiwamu]

 We have installed the PRM and the tip-tilt (TT) in the BS chamber.

We have started the in-vac work which takes about a week.

Today's mission was dedicated to installing the PRM and two TTs, one for the PRC and the other for the SRC, on the BS table in the chamber.

The work has been smoothly performed and we succeeded in installation of the PRM and a TT for the PRC.

But unfortunately the other TT got broken during its transportation from Bob's clean room.

 


 (what we did)

 (1) opened the light door of the BS chamber.

 (2) moved the BS tower to the right position according to Koji's layout drawing.

    - Prior to this work we screwed down the earthquake stops so that the mirror is fixed to the tower. Also we disabled the watchdog.

    - When moving it we used an allen key as a lever with an screw as a fulcrum. This idea was suggested by Jenne and it really worked well.

     The reason why we used this technique is that if we slide the tower by hands the tower can't go smoothly and it may sometimes skips.

     After that we checked the postion from some reference screw holes by using a caliper and we made sure that it was on the right position.

 (3) removed all of the square-shaped mirrors.

    - After this removal the mirrors were wrapped by aluminum foils and put in a usual clear box.

 (4) removed some optics because they had made the chamber space crowded.

    - These were also wrapped by aluminum foils and put in the box. Later we will put them back to the BS table.

 (5) brought the PRM tower from the Bob's clean room  and put it on the BS table.

   - The position of the PRM were coarsely aligned since we still don't have any 1064 beam going through the PRM.

 (6) brought two TTs also from Bob's clean room and put one of the TTs on the table.  

   - The position of the installed TT was coarsely adjusted. 

   - After we brought them we removed the aluminum foils covering the TTs and we found the wire of a TT got broken.

     It may have been damaged during its transportation from Bob's room because it was fine before the transportation.

 (7) closed the door

 

(the next things to do)

  * Installation of the OSEMs to the PRM

  * Installation of the pick off mirror and its associated optics

  * Arrangement of  the pzt mirror

 

Attachment 1: DSC_2275.JPG
DSC_2275.JPG
  3277   Fri Jul 23 15:32:02 2010 steveUpdatePEMcontrol room AC set point changed

The control room temp is warmer than usual. The heat exchanger Office Pro 18 set point was lowered from 70 to 68F yesterday.

The MOPA headtemp is higher also. The Neslab chiller bath temp peaks around 21.6 C daily. This should be rock solid 20.00 C

It did not have any effect.

Now, I have just lowered the thermostat setting of room 101 from 73 to71F  I  hope Koji can take this.

  3276   Fri Jul 23 14:26:01 2010 GopalUpdateOptic StacksSimple Frequency Response Measurements in COMSOL

Over the past couple days, I discovered a simple, direct method for calculating frequency responses with a combination of COMSOL and any plotter such as Excel or MatLab. The simple case of rectangular prism of steel was analyzed using this method; details will be posted shortly on the COMSOL Wiki page. The frequency response matched theoretical reasoning: the bar acts as a simple mechanical low-pass filter, rapidly attenuating driving frequencies at the base beyond the first eigenmode.

It therefore shouldn't be too difficult to extend this analysis to the MC1/MC3 stack. The many eigenfrequencies will produce a more complicated transfer function, and so more data points will be taken.

The major shortcoming of this method involves dealing with the imaginary components of the eigenfrequencies. As of now, I haven't found a way of measuring the phase lag between the drive and the response. I also haven't found a way of changing the damping constants and therefore playing with phase components.

 

  3275   Fri Jul 23 12:33:51 2010 josephbUpdateCDSmegatron, c1iscex, c1sus firewalled

Kiwamu and I strung an ethernet cable from the new 1X7 rack to the 1X3 rack. The cable is labeled c1iscex-daq on both ends.  This cable will eventually connect c1iscex's second ethernet port to the daq router.  However, for today, it plugged into the primary ethernet port and is going to a linksys router.  This is the same linksys router we used to firewall megatron.

The idea is to place megatron, c1sus, and c1iscex behind the firewall to prevent any problems with the currently running while doing RFM nework tests.

The way to get into the firewalled sub-network is to ssh into megatron.  The router will forward the ssh to megatron.  Inside the network, the computers will have the following IPs.  Router is 192.168.1.1, megatron is 192.168.1.2, c1sus is 192.168.1.3, and c1iscex is 192.168.1.4.

 

 

  3274   Fri Jul 23 10:16:44 2010 josephbUpdateCDSRerouted RFM around c1lsc, took RFM card out of c1lsc

 I re-routed around the c1lsc machine this morning.  I turned the crate off, and disconnected the transmission fiber from  c1lsc (which went to the receiver on c1asc).  I then took the receiving fiber from c1lsc and plugged it into the receiver on c1asc. 

I pulled out the c1lsc computer from the VME crate and pulled out the RFM card, which I needed for the CDS upgrade.  I then replaced the lsc card back in the crate and turned it back on.  Since there hasn't been a working version of the LSC code on linux1 since I overwrote it with the new CDS lsc code, this shouldn't have any significant impact on the interferometer.

I've confirmed that the RFM network seems to be in a good state (the only red lights on the RFM timing and status medm screen are LSC, ASC, and ETMX).  Fast channels can still be seen with dataviewer and fb40m appears to still be happy.

The RFM card has found its new home in the SUS IO Chassis.  The short fiber that used to go between c1asc and c1lsc is now on the top shelf of the new 1X3 rack.

  3273   Fri Jul 23 08:18:03 2010 josephbUpdateCDSNot enough room in IO chassis for RFM card - need to swap PMC to PCIe adapters

The End IO chassis have small trenton boards, which apparently only have 5 usuable PCI slots, even though there are 6 on the board.  This is because of the way the the host interface board is setup and its closeness to the 2nd to last PCI slot

The PMC to PCIe adapters I was handed by Jay for use with the RFM cards require a 4 pin power connection at the top, which are not available inside the thin 1U computers.

The only solution I can come up with is swap the PMC to PCIe adapters for the RFM cards with adapters for some of the already installed ADCs and DACs which do not require power directly from the power supply.  This should make it possible to mount the RFM card in the computer, at least for the ends.  Since the SUS and IOO chassis will have more slots available than needed, the RFM cards can be slotted into those.  The SUS has to fit in the chassis since the computer will have the Infiband host adapter and a dolphin connector for talking to the LSC machine.

There is still the problem of actually getting the RFM card into the computer, but that should be possible with a little bit of bending of the left side of the computer frame.

  3272   Fri Jul 23 08:15:59 2010 steveUpdatePSLref cavity ion pump

The ref cavity ion pump was running at 7.7kV instead of 5kV

This Digitel SPC-1 20 l/s ion pump should be running at 5kV

  3271   Fri Jul 23 00:13:11 2010 ranaUpdatePSLProblem NOT REALLY Solved

So...who was working around the PSL rack this morning and afternoon? Looks like there was some VCO phase noise work at the bottom of

the rack as well as some disconnecting of the Guralp cables from that rack. Who did which when and who needs to be punished?

  3270   Thu Jul 22 18:18:54 2010 AlbertoUpdatePSLProblem Solved

Quote:

Quote:

It looks like something wrong happened around the PSL front end.  One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy. 

We found it by the donkey alarm 10 minutes ago.

The attached picture is a screen shot of the PMC medm screen.

The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.

Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)

The problem seems to be a software one.

In any case, Kiwamu and I looked at the at the PMC crystal board and demod board, in search of a possible bad connection. We found a weak connection of the RG cable going into the PD input of the demod board. The cable was bent and almost broken.

I replaced the SMA connector of the cable with a new one that I soldered in situ. Then I made sure that the connection was good and didn't have any short due to the soldering.

[Alberto, Koji]

By looking at the reference pictures of the rack in the wiki, it turned out that the Sorensen which provides the 10V to the 1Y1 rack was on halt (red light on). It had been like that since 1.30pm today. It might have probably got disabled by a short somewhere or inadvertently by someone working nearby it.

Turning it off and on reset it. The crazy LO calibrated amplitude on the PMC screen got fixed.

Then it was again possible to lock PMC and FSS.

We also had to burtrestore the PSL computer becasue of the several reboots done on it today.

  3269   Thu Jul 22 15:59:29 2010 AlbertoUpdatePSLPSL front end machine

Quote:

It looks like something wrong happened around the PSL front end.  One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy. 

We found it by the donkey alarm 10 minutes ago.

The attached picture is a screen shot of the PMC medm screen.

The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.

Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)

The problem seems to be a software one.

In any case, Kiwamu and I looked at the at the PMC crystal board and demod board, in search of a possible bad connection. We found a weak connection of the RG cable going into the PD input of the demod board. The cable was bent and almost broken.

I replaced the SMA connector of the cable with a new one that I soldered in situ. Then I made sure that the connection was good and didn't have any short due to the soldering.

  3268   Thu Jul 22 14:07:20 2010 kiwamuUpdatePSLPSL front end machine

It looks like something wrong happened around the PSL front end.  One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy. 

We found it by the donkey alarm 10 minutes ago.

The attached picture is a screen shot of the PMC medm screen.

The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.

Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)

Attachment 1: PSL-PMC2010-07-22.png
PSL-PMC2010-07-22.png
  3267   Thu Jul 22 13:44:47 2010 JennaUpdatePEMGuralp seismometer

One of the Guralps [Gur2] has been taken to the atf gyro lab, along with the breakout box.

 

Edit by Jenne:  This means that we have no working seismometers in the 40m lab right now, so don't worry if you're looking for seismo data and you can't find any.  The 6 accelerometers should all still be up and running.

  3266   Thu Jul 22 08:07:27 2010 steveUpdateVACvent finished

Quote:

The vent has been finished successfully in this morning.

The vent was finished successfully this morning.

 Thanks to Kiwamu, Alberto and Koji

Attachment 1: 30hvent.jpg
30hvent.jpg
  3265   Thu Jul 22 07:19:56 2010 AlbertoUpdateTreasureMonsters, LNVR, and Phase noise

Quote:

On Picasa

 "They (shellfish) shall be an abomination to you; you shall not eat their flesh, but you shall regard their carcasses as an abomination." (Leviticus 11:11)

  3264   Thu Jul 22 03:08:27 2010 KojiUpdateIOOVent and MC lock

Summary

- The vacuum chambers have been vented.

- The north heavy door of the BS chamber has been opened by Genie (not by the crane).
It was replaced by the light door. The door is currently closed.

- The MC has been locked with 20mW incident and aligned. MC REFL was left unchanged but lock was able to be achieved.

- The optics before MCT CCD and MCT QPD have been adjusted for the low power operation.


Details

- The first HWP for the variable optical attanuator (HWP/PBS/HWP pair) was set to be 86deg from the maximum transmission at 126deg.
The incident power of 19mW has been measured.

- The PSL mechanical shutter has been manually opened. Two other beam blocks has been removed.

- I found the MC was totaly misaligned with no resonance.- I tried to align it based on the previous OSEM values but in vain.

How to align the MC mirrors from the scratch

- MC1 has been aligned so as to maximize REFL PD and DC signal of WFS QPDs.

- MC3 has been aligned by looking at the scattered light on the MC2 frames. The spot is centered on the MC2 approximately.

- MC2 has been aligned so that any resonance is seen in MC_F.

Modification of the MCT optics

- The ND filter before the MCT was removed.

- The Y1-45S mirror before the MCT CCD, which is also used to steer the beam to the MCT QPD path, was replaced to BS1-50-45P.
The reason I used 45P is to obtain higher reflectivity. Because S has higher reflectivity than P in the each layer, I expected to have higher reflectivity for S than 50%.

- The MC REFL path has not been untouched.

Modification of the servo

- The lock was attempted after alignment of the mirrors.  Here how to lock the MC is described below.

1. Run script/MC/mcloopson

2. Open the MC Servo screen in MEDM

3. Change the input gain from 6dB to 22dB.
Change offset from 0.78 to -0.464 (such that the length output has no offset).
Change VCO Gain from 3dB to 21dB

Change MC Length path Gain from 0.3 to 1.6

  3263   Thu Jul 22 01:02:08 2010 ranaUpdateTreasureMonsters, LNVR, and Phase noise

On Picasa

  3262   Wed Jul 21 19:11:18 2010 DmassUpdateGreen Lockinglocked

What did you use to filter the 2f components from your error signal? A homemade low pass or what?

 


Kiwamu:

I am using a homemade low pass filter.

It's just a RC passive LPF with the input impedance of 50 Ohm.

  3261   Wed Jul 21 17:41:17 2010 GopalConfigurationOptic StacksPictures of Stacks

Now that venting is complete, this is a request for anyone who opens any chamber:

1) Please notify me immediately so I can take pictures of the stacks in that chamber.

2) If I'm not around, please take a few pictures for me. I'm most interested in the shape, number of layers, size, and damper arrangements of each stack.

This is most important for the MC1/MC3 chamber, MC2 chamber, and BS/ITMX/ITMY chambers.

Thanks!

ELOG V3.1.3-